A Complete Bibliography of Publications in the ACM Symposia on Theory of Computing (STOC) for

1970–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 September 2023 Version 1.03

Title word cross-reference

2.5n [Pau75]. d [Ber79]. γ [Lon79]. L [ERS78]. n [LD75]. n log n [Mil78]. O(v^{O(g)}) [FMR79]. Ω(n^{2} log n) [YAR77]. p [MY76].

-axiomatizable [Ber79]. -hierarchy [MY76]. -linear [LD75]. -lower [Pau75]. -reducibility [Lon79].

Aanderaa [RV75]. Aanderaa-Rosenberg [RV75]. Absolutely [Raj71]. Abstract [AM77, Ang79, Bea77, Ber71, BGW70, BNP74, Bur76, CW77, ER75a, ER75b, ERS78, GHR76, Gre73, GS76, Ham74, HS76, HPV74, KF77, Lad79, LRS73, Lon79, Lue78, Mai79, PF77, Ros70b, Ros71, Ros74, RS75, Ruz79, Sav72, SS78b, cY76, Ham72, LR70, Meh74, Ull70]. abstractions [Doe76]. accepted [Coo79]. acceptors [BGW70]. access [CR72, FW78]. ACM [ACM75, ACM76, ACM77, ACM78]. addition [BM78, ST77, vL74]. additions [BC74, Kir72]. Addressable [Ros71]. addressing [Ros70b]. admissible [ST78]. AFDLs [Gre73]. AFL [CGGG72, GG71]. AFLs [BGW70]. Albuquerque [ACM75]. algebra [Val79a]. Algebraic [San71, Kim76, Kun73, Tei73, Wag71]. algebras [Koz77, Wag73]. algorithm [Ano78, Fid72, Fil78, HW74, HPV74, Pag73, Ter71, Wis71]. Algorithmic [RT75]. Algorithms [GK78, Hor72, AV77, BL75, Gal76, GGU72, GJ73, Hir76, Joh73, Kos79a, Kun74, Mun71, Rac78, WSB73, cY76]. allocation [GGU72, Set73]. allocations [Set72]. alpha [Ano78]. alpha-beta [Ano78]. alternation [RL78, Ruz79]. Analysis [GJ73, Kos73, Ano78, Cha70, GGU72, GM77, GS76, Pra75, Tei73]. ancestors [AHU73]. Annual [ACM72, ACM73, ACM75, ACM70, ACM71, ACM74, ACM76, ACM77, ACM78, ACM79]. Anomaly [CS78]. application [FFV79, Wan73]. applications [Fre75, LP76]. applied [HM73, Wis71]. approach [Gol78]. approximate [CFG^{+}78]. Approximation [Joh73]. April [ACM73, ACM74, ACM79]. Area [Tho79]. Area-time [Tho79]. arguments [Ked74]. arithmetic [Dow78, Hor72, Opp73, RL78, O’D79]. array [Kos79a]. arrays [KMR72, Ros74, RS75, Sei74]. aspects [RT75]. assertions [OC75]. Associative [Bur76]. asynchronous [Pet79]. Atlanta [ACM79]. Attributed [LRS73]. Austin [ACM73]. Author [Ano77]. automata [But71, LJ71, Per70, SS78a, Sav72, Sip79, Sud76, Sud77, Val74, Ell70]. auxiliary [Sud76, Sud77]. average [cY76]. axiomatic [HPS77]. axiomatics [CV74]. axiomatizable [Ber79].

balance [NR72]. balanced [Fre75]. based [BY70, Sch76]. bases [CM77, CH79]. be [Val79b, WSB73]. behavior [cY76]. beta [Ano78]. between [DL79, Har77, HR74, LB77, Mor70, Per70, RV78]. bilinear [BD73, HM73, Ja’78, Ja’79]. Binary [NR72, DL74, RMK78]. bipartite [GK78]. Block [Ber71]. Boolean [FMP75, Pau75, Pip76, Pra74]. bottom [Ter71]. Boulder [ACM77]. bound [Kun72, Pau75, YAR77]. bounded [BGW70, BNP74, CR72, Gal75, NR72, RL78, Ruz79, Sud77]. bounds [Aus70, BFP^{+}72, FMP75, Ked74, Kun74, LT79, Opp73, PTC76, Sip79, Val75]. branching [Lip74]. building [Fre75].

calculus [CR74, Coo75]. California

[ACM78]. can [Val79b]. canonical [Cha74]. care [AU71]. case [GGU72, GJ73]. certain [Kir72, Kun74]. CFL [Coo79]. characterization [PRS74]. Characterizations [WS72]. charts [Str70]. chosen [Lue78]. Church [Ros70a]. circuit [IR77]. circuits [AV77, Tom78, Val76]. class [BY70]. Classes [Eil73, BM78, CW77, Koz78, LR70, Mac71, Meh74, Rob71, Sud77, Val79a]. Classification [Mac71]. closed [GG71]. Closure [Lew70a, Hir76, Mil74a]. code [AJ75]. codes [Per70]. coding [Bur76]. Colorado [ACM72, ACM77]. coloring [GK78]. combinational [Pau75]. Combinatorial [Meg78, ER75a, ET75, Joh73]. combinatory [O’D77]. Combining [Ked74]. common [AHU73, Ang79]. communicating [Lad79]. communication [Kim76]. commutative [CLM76]. commutativity [Ja’79]. Comparative [GL75]. Comparison [LLS74, PS77]. Complete [JL74, Set73, CLM76, ET75, GJS74, GGJ76, GLT79, GI78, HPS77, HB76, MA76, Owi76, Yan78]. Completeness [Val79a, Ber79, HMP77, KH78]. complex [Pan78]. Complexity [Bur70, CH71, EZ74, Koz77, LD75, Rob71, Sch76, BY70, Bro77, Con73, Coo71, Coo72, DL79, Gal75, Gil74, GL75, GI79, Har77, HM73, Hun73, HS75, HK75, Ja’79, Kun73, Lad79, LR70, Lew70b, Lew78, LED75, Lyn78, Mil74b, Pan78, PS76, Pau75, Rac74, Rob74, Sch78, Sch72, SS79, Sha75, Tho79, Val75, Wag75, Weg78, Yao79]. component [Hir76]. Compositions [OR72]. compression [GN79, SS78b].

Computability [EC75, HMP77, ST78]. Computable [CH79, Mac71]. computation [Bur70, Cha71, Man70, Moe73, Sym72, VD74]. Computational [Gil74, HR74, Mil74b, Pan78, BY70, Con73, Kun73, Lew70b, Val75]. Computations [KF77, BFP^{+}72, Ham72]. compute [Aus70, BC74, Kir72]. computer [TK76]. Computing [ACM70, ACM71, ACM72, ACM73, ACM74, ACM75, ACM76, ACM77, ACM78, ACM79, FFV79, Pra73, HPV74, Pan78, Tom78, Yao75, Yao79]. concatenable [Kos79c]. concepts [EC75]. concerning [Mun71]. conditional [Lip74]. Conference [ACM70, ACM71, ACM72, ACM73, ACM74, ACM75, ACM76, ACM77, ACM78, ACM79]. conjecture [RV75]. conjunctive [CM77]. connected [Hir76, TK76]. connections [DL79]. connectivity [Tar74, Tom78]. conquer [BS76]. consecutive [BL75]. considerations [Lew70b]. consistent [Owi76]. constructing [Pag73]. Construction [Mil74a]. constructive [Coo75]. Context [Tei73, BNP74, GHR76, Gre73, HR74, Rou70, Sud76, Ter71]. Context-free [Tei73, BNP74, Gre73, HR74, Rou70, Sud76]. control [LED75, Pra75]. convex [CD79]. Coping [RMK78]. Correct [Vui73]. Correction [Wag75]. correctness [McG70]. costs [FFV79]. counter [Boa71, GI79]. counters [Pla72]. covers [Mai79]. critical [Kat78, PF77]. cryptocomplexity [Sha79]. cubic [Fil78, GHR76].

Data [Ros70b, TWW78, CM77, CH79, Fre79, LED75, MS79, OC75, Ros71, RWG79, SS78b]. database [BMSU79, Mai79]. Deadlock [TU79]. Deadlock-free [TU79]. decidability [ST77, Val74]. Decidable [CM72b]. decision [IK76, MA76, McN71, Pra78, Sch76, Ter71].

Decomposing [CD79, Sky75]. deductions [HPS77]. deductive [Owi76]. defined [Sky75]. definitions [Wag71]. Degree [Wot75, Cud70]. Degree-languages [Wot75]. Degrees [Cha74]. deletion [Ber71, Yan78]. density [HB76]. Denver [ACM72]. depth [FU76]. derivatives [Mil74a]. descriptions [MB72]. design [Lin70]. Determinants [GJ73]. Determining [Kir74, BY70, Fil78, FMR79]. Deterministic [Coo79, CS74, Gre73, JL74, Raj71, Sud76, Ter71, Val74]. deterministic/polynomial [CS74]. Developmental [LR72]. device [IK76]. devices [Sky75]. diagonalization [Har77]. diagonalizations [SG77]. Dichotomization [HS76]. dictionaries [FFV79]. Diego [ACM78]. digraphs [VTL79]. dimensional [Cha70]. dimensionality [Ked74]. direct [Rac74]. directed [MV70]. disjoint [Tar77]. disk [HPV74]. distributed [PF77]. distribution [Lue78]. distributive [Yao79]. Divide [BS76]. Divide-and-conquer [BS76]. division [Fid72]. domains [ST78]. Domolki [Wis71]. don’t [LMF73]. double [GS76, Kos79c, Pra73]. double-ended [Kos79c, Pra73]. Draft [MS79]. Duality [HM73]. Dyck [Niv70]. dynamic [Pra78, Ros77, MW79].

Economical [PF77]. economy [MB72]. edge [GK78, Lue78, Yan78]. edge-deletion [Yan78]. effective [Sch72]. efficiency [Kun72]. Efficient [Hor74, LB77, Fil78, HPV74, Mun71]. eighth [ACM76]. Elementary [Opp73]. eleventh [ACM79]. elimination [Rei70, RT75]. ended [Kos79c, Pra73]. enumerable [GG71]. equality [CM72a, JS72]. equations [O’D77, Str70]. Equivalence [BMSU79, CM72b, GL72, GI79, RV78, Val74]. Equivalences [Pet79]. erasable [CGGG72]. error [Mil73, Tei73]. errors [RMK78]. establishing [Ked74]. Evaluation [LS76, BD73, Fid72, HK75, Hya78, Ja’78, Kun74, LD75, Tei73]. Exact [CFG^{+}78]. Exponential [CLM76, SM73]. exponentially [Val79b]. expression [AJ75]. expressions [EZ74, Kun74]. expressive [MW79]. Extended [Ang79, Bea77, Ber71, BGW70, BNP74, Bur76, CW77, ER75a, ER75b, ERS78, GHR76, Gre73, GS76, Ham74, HS76, HPV74, KF77, Lad79, LRS73, Lon79, Lue78, Mai79, PF77, Ros70b, Ros71, Ros74, RS75, Ruz79, Sav72, SS78b, cY76, Wag75]. extendible [Ros74, RS75]. extra [Hor74].

families [Bak73, Lew70a, Niv70, Ull70]. Fast [AV77, Fid71, FS73, Kos79a, Moe73, Fid72]. feasible [Sim77]. Feasibly [Coo75]. feeding [AU71]. fields [Pan78]. Fifth [ACM73]. Finding [Ang79, FU76, IR77, AHU73, SS79, WSB73]. fingers [BT78]. finite [But71, Per70, Raj71, SS78a, Sch76, Sym72, Val74]. finite-state [Raj71]. finite-turn [Val74]. finitely [Koz77]. first [JS72]. first-order [JS72]. fixedpoint [MS75]. Flow [HU72, AU75, FU76, GN79, Str70, Tar73]. Flowchart [Pla72]. flowchartable [WS72]. flowcharts [Kos76]. forbidden [HS76]. form [Ham74]. Formal [Sta71, CH71, Cud70, ER75b]. formalisms [Con70]. forms [BD73, Cha74, GL75, HM73, Ja’78, Ja’79, LD75, Yao75]. formulas [FMP75, JS72]. formulating [MY78]. Four [Pra75]. Fourier [Fid72]. Fourth [ACM72]. free [BNP74, GHR76, Gre73, HR74, Mey71, Rou70, Sud76, Tei73, Ter71, TU79]. full [Ano78, GG71]. function [Hor72]. functions [Aus70, CW77, Kir72, Mac71, Meg78, Pau75, Pip76, Rob71, Sym72, Tom78, WF72, Weg78]. game [PTC76]. games [Rei79, Sch76]. gaps [Har77]. GCDs [Moe73]. general [Mil77]. generalization [RV75]. generalizations [DL74]. generalized [But71, GN79, KH78, McN71, Wis71]. generating [Ter71]. generation [AJ75]. generative [But71]. genus [FMR79]. Geometric [Sha75, GGJ76]. Georgia [ACM79]. global [Lip74]. Gödel [MY76]. grammar [GL75, HS75]. grammars [AU71, But71, Cha70, Cud70, Gre74, Raj71, Ter71, Wis71]. grammatical [Ham74]. Graph [Mil77, Fil78, FMR79, FU76, HU72, IR77, Kir74, Kos79a, Tar73, Tar74]. graphs [AU75, BL75, GK78, HW74, Lue78, PTC76, Ros70b, Ros71]. growth [Ked74].

Hamiltonian [AV77]. hash [Bur76, CW77]. hash-coding [Bur76]. Hashing [RS75, GM77, GS76]. Heights [ACM71]. help [LMF73]. Hershey [ACM76]. Hierarchies [BY70, CS78, LD75, Pau77]. hierarchy [Coo72, Cud70, MY76]. Hoare [Wan76]. homeomorphism [LR78]. Hypothesis [Mil75].

identification [KMR72]. II [CM72b].

implementation [CM77].

implementations [Vui73]. Implicit [MS79]. improved [GM77]. incomplete [Rei79]. incompleteness [Wan76]. incremental [Lin70]. independent [Ham72, O’D79]. index [Ano77]. indexed [Rou70]. Indexing [Koz78]. inductive [CS78]. inefficient [WSB73]. inference [CS78]. information [Rei79, Sch76]. instruction [PS77]. integer [FS73]. integers [KF70, LD75]. integrated [FFV79]. interactions [ER75a]. Intercalation [Per75]. interpretation [KF70]. interprocess [Kim76]. Intersection [GG71]. Intersection-closed [GG71]. Intersections [BNP74]. interval [BL75]. intractibility [AM77]. involving [Rac78]. irregularities [BY70]. isomorphism [HW74, Mil77, Mil78]. isomorphisms [HB76]. Iteration [Bea77, Boa71, Kun72]. iterative [Sei74].

Jump [Gre73].

knapsack [Sha79].

L [ER75a]. lane [Pag73]. language [Coo70, CV74, EC75, GHR76, Lin70, Niv70, O’D79, Vui73]. languages [Bak73, Bea77, Boa71, BNP74, CS74, ER75a, ER75b, GG71, Gre73, Ham72, Hun73, HR74, Lew70a, LR72, Mor70, Niv70, Per75, Rou70, Sky75, Sta71, Sud76, Ull70, Wag71, Wag73, Wot75]. LBA [Wot75]. length [Lyn78]. lengths [CR74]. less [GHR76]. like [RWG79]. Limitations [Lip74]. line [AU70, FS73, GHR76, Lyn78, Set72]. Linear [BFP^{+}72, BL75, HW74, Mey71, PW76, BNP74, BT78, GMPR77, HK75, LD75, Tar77, Val75]. LISP [O’D77]. listings [AU75]. lists [BT78, GMPR77, Mor72]. LL [Bea77, Ham74]. local [Cha71]. Location [LP76]. log [Coo79]. logic [DL79, FL77, O’D77, Pra78, MW79]. logics [Con77, HMP77]. Loop [Con71]. Lower [FMP75, Sip79, Ked74, Kun74, LT79, Pau75, Val75, YAR77]. lowest [AHU73]. LR [AU71, AU72, Pag73]. lucid [O’D77].

machine [Ham72, McG70]. machine-independent [Ham72]. machines [BNP74, CR72, ERS78, FW78, Gil74, Gol78, GI79, JS72, PS77, PRS74, San71, Tar77, WF72]. macro [SS78b]. maintain [Tar77]. maintenance [Bro77]. Managing [Ros74]. manipulate [OC75]. manipulating [Ros70a]. many [Lon79]. many-one [Lon79]. mapability [McN71]. mapability-onto [McN71]. Massachusetts [ACM70]. matching [Gal76, KH78]. matchings [AV77]. mathematical [DL79, Man70, Mey71]. matrices [Pra74]. matrix [Fid71, HM73, Kir74]. Maximization [Lue78]. Maximum [Lew78]. May [ACM70, ACM71, ACM72, ACM73, ACM74, ACM75, ACM76, ACM77, ACM78, ACM79]. Maze [Sav72]. measured [BY70]. measures [EZ74, LD75, Lyn78]. mechanical [Mil73]. mechanisms [HA77]. mechanized [CS78]. median [BFP^{+}72]. membership [CFG^{+}78]. memory [GGU72]. merge [HPV74]. merging [cY76]. mesh [TK76]. mesh-connected [TK76]. method [Pra78]. Mexico [ACM75]. minima [Yao75]. minimal [Hor74]. Minimum [Mai79, IR77]. modal [FL77]. model [Mai79, SS78b, VD74]. models [Gol78, Pet79, Pra75]. modified [McG70]. monadic [Rob74]. monoids [Per70]. monotone [Pip76, Weg78]. movable [BT78]. multidimensional [BS76, Kos79b, Sei74]. multihead [LS77, Sud76]. multiplication [Fid71, FS73, Kun72]. multiplications [HM73, Ked74]. multiplying [Pra74]. multipushdown [BNP74]. multivariate [Hya78]. must [WSB73].

n [OR72]. natural [Rob74]. near [Fre79]. necessary [Kir72]. Negation [Val79b]. negative [Pra74]. Neighborhood [WSB73]. nets [CLM76]. Network [GN79]. networks [TU79]. ninth [ACM77]. Node [AU75, Yan78]. Node-and [Yan78]. non [Tar77, Val75]. non-linear [Tar77, Val75]. Nondeterminism [SS78a, HA77]. nondeterministic [Coo72, KF77, Sei74]. Nonserial [Ros77]. normal [Lue78]. Northampton [ACM70]. notations [Mey71]. NP [GJS74, GGJ76, GI78, HB76, MA76, Yan78]. NP-complete [GJS74, GGJ76, GI78, MA76, Yan78]. number [Aus70, BC74, GI78, Ked74, KF77]. number-theoretic [GI78]. numberings [MY76]. numbers [Kun73, Pan78, Rob74, Sim77]. numerical [Mil74b].

objective [Meg78]. observation [Coo73]. Observations [Sei74]. off [Coo73]. offs [Pet79]. Ohio [ACM71]. on-line [FS73]. one [Boa71, Lon79]. one-counter [Boa71]. ones [BL75]. onto [McN71]. Operational [RV78]. operations [Cha71, FFV79, Hor72]. operators [Lew70a]. Optimal [AJ75, CM77, Ja’78, BD73, Fre79, HPV74, MS75, Mun71, Ros77, Vui73, WSB73]. optimization [Meg78, Pra75]. order [JS72, Man70, Rob74]. oriented [KF70, Rou70]. other [HB76, HM73]. overlap [Wis71].

packet [TU79]. pairs [Ja’78]. palindrome [Gal76]. papers [ACM70, ACM71, ACM72, ACM73, ACM74, ACM75, ACM76, ACM77, ACM78, ACM79]. Parallel [Hir76, VD74, Cha71, Doe76, Gol78, HK75, Hya78, Kos79a, Kun74, Mil74a, Owi76, Pet79, Pra73, Raj71, Sky75, TK76, VTL79]. Parallelism [FW78, MY78]. parallels [HR74]. Parameter [HA77, Lyn78]. Parameter-passing [HA77]. Parameterization [TWW78]. parenthesis [Mey71]. parenthesis-free [Mey71]. parser [Ter71]. parsers [AU72, Lin70, Pag73]. Part [Cha74]. partial [Rob71, vL74]. parts [CD79]. passing [HA77]. patent [Ham72]. Path [Coo70, GN79]. paths [YAR77]. patterns [Ang79, Cha70, HPV74, KMR72]. PDA [Gre73]. Peano [O’D79]. pebbling [GLT79]. Pennsylvania [ACM76]. perfect [Sch76]. perfect-information [Sch76]. periods [SS79]. permutations [Pra73]. person [Sch76]. Petri [CLM76]. picture [Cha70]. planar [HW74, LP76]. point [LP76]. polygon [CD79]. Polynomial [Fid72, Meh74, SG77, CS74, Coo79, ET75, GLT79, Gre73, JL74, LLS74, Lon79, Wot75]. polynomial-time [LLS74]. polynomials [BC74, GJ73, Hya78, LD75, LS76, MA76, Pan78]. Polynominal [Lad73]. power [MW79, PRS74, Pra74, Sta71, Sud76, Tei73, TWW78]. powerful [Val79b]. practical [Pra78]. precedence [Mor70]. preconditioning [LS76]. Predecessor [WF72]. predicate [Rei70]. predictability [ER75b]. prefix [Per70]. Preliminary [AU70, BY70, BC74, CLM76, CH79, CR74, Coo75, Dow78, FMR79, HMP77, HW74, Kos76, Kos79a, Kos79b, Kos79c, LR70, LB77, MY76, MW79, Mil78, MS79, Pip76, Pra78, Rac74, RMK78, Rob71, ST77, Sim77, SM73, Val76, Yao75, Yao79, FMP75, MB72]. Presburger [Opp73, RL78]. presented [ACM70, ACM71, ACM72, ACM73, ACM74, ACM75, ACM76, ACM77, ACM78, ACM79, Koz77]. primality [Mil75]. primitive [Mac71]. primitives [Lip74]. principal [Gre73]. priority [Bro77]. Probabilistic [Ell70, AV77, Fre75, Gil74, Rac78]. problem [ET75, Fre79, GLT79, GI78, GI79, Ham72, HS76, Kat78, KH78, Kos79b, LR78, PF77, ST77, Wot75, YAR77, vL74, Lew78, PS76, Wag75]. problems [Bur70, CLM76, CM72b, Cud70, GJS74, GGJ76, Hir76, HS75, IK76, Joh73, JL74, Kos79a, Lad79, Lue78, MA76, Sch76, Sch78, Set73, SM73, Yan78]. procedure [McN71]. procedures [Coo71, RMK78]. process [Kim76, Sch72]. processes [Lad79]. processing [Cha70, Kos79a]. processors [Lin70]. products [Rac74]. Program [CM72a, MB72, Cha74, CM72b, Lyn78, Mil74a, Pra75, Wan73]. program-proving [Wan73]. programming [Con77, EC75, Ham72, LB77, O’D79, Ros77, Vui73, Wag73]. programs [AU70, Con70, Doe76, FL77, GI79, HPS77, HMP77, Kos73, MS75, OC75, Owi76, RV78, Set72, VD74]. proof [Har77, RV75]. proofs [CR74, Coo75, Dow78, Rou70]. propagation [Mil73]. properties [Kir74, LR70, Mil73, Tom78]. property [BL75]. Propositional [Dow78, FL77, CR74, Coo75, Pra78]. prover [KF70]. Proving [OC75, Coo71, HPS77, Rei70, Wan73]. pruning [Ano78]. pushdown [Sud76, Sud77, Val74].

quadratic [MA76, Yao75]. quantifier [RL78]. queries [CM77, CH79]. query [Fre79]. questions [Rac78, Yao79]. queue [Bro77]. queues [Kos79c, Pra73].

random [CR72, FW78, Sch72]. randomness [AM77]. range [Fre79]. Rapid [KMR72]. rate [Ked74]. rational [Hor72, Kun74, Meg78]. reachability [HS76, ST77, vL74]. reachability-problem [vL74]. Real [Gal76, Kos79c, BM78, Bur70, LS77, Pan78]. Real-time [Gal76, Kos79c, LS77]. realization [Pip76]. recognition [Coo70, Gal76, GHR76, Gre73, VTL79, Wot75]. recognizable [CS74]. recognize [BL75]. recognizing [Sav72]. record [ACM70, ACM71, ACM72, ACM73, ACM74, ACM75, ACM76, ACM77, ACM78, ACM79]. recurrence [HK75]. Recursion [Mor72, Str70, Vui73]. recursions [WS72]. Recursive [LR70, CV74, HPS77, Mac71, MS75, O’D77, RV78, Rob71, Wag71]. recursively [GG71]. Reducibilities [Pet79, LLS74, SG77]. Reducibility [AM77, HU72, Lad73, Lon79, LB77, Tar73]. reducible [AU75]. reducing [Mor70]. Reference [Tar77]. register [Set72, Set73]. regressing [WF72]. regular [EZ74, HR74, McN71]. related [HS75, Niv70, Yao79]. relation [BM78]. relational [BMSU79, CM77, CH79, Mai79]. Relations [Har77]. relationship [Mor70, Per70]. Relativized [Rac78]. remarks [Mil77]. repeated [KMR72]. replacement [O’D77]. Report [BY70, CLM76, CH79, FMR79, HMP77, HW74, LB77, MY76, MW79, Mil78, Pra78, Rac74, RMK78, SM73, Val76, Yao75, Yao79, FMP75, MB72]. representation [BT78, Dow78, GMPR77, Mey71]. representations [Kir74, RWG79]. require [Tar77]. requirements [CS74]. requiring [SM73]. resolution [Gal75]. resolvable [Wis71]. restricted [KF77, VD74]. restrictions [Gre74]. result [Mor70, Wan76]. results [CH71, LJ71, Mun71, PS76]. Retention [Ber71]. retrieval [Bur76]. reversal [BNP74]. reversal-bounded [BNP74]. revisited [Fid72]. Riemann [Mil75]. Rosenberg [RV75]. Rosser [Ros70a]. roundoff [Mil73].

Salesman [PS76, WSB73]. San [ACM78]. satisfiability [Sch78]. scheduling [VD74]. schema [Mil74a]. schemas [CM72a, Cha74]. schemata [Con71, CM72b, Pla72]. schemes [BMSU79, GL72, Mor72, Ros70b, RS75]. search [DL74, Fre75, Kos79b, NR72, RMK78, WSB73]. Seattle [ACM74]. SECD [McG70]. Second [Man70, ACM70, Rob74]. Second-order [Man70, Rob74]. section [Kat78, PF77]. semantic [RV78]. Semantics [CV74, Ber79, EC75]. semigroups [CLM76, Eil73]. semilinear [GI79]. Separating [Sud77]. sequences [FFV79]. sequential [Lad79, McN71]. series [Sta71, Tei73, VTL79]. set [Ang79, BD73, cY76]. Sets [LMF73, BY70, Eil73, GI79, HB76, McN71, PS77, Tar77]. Seventh [ACM75]. Shaker [ACM71]. short [WS72]. shortest [YAR77]. showing [IK76]. Simple [MY76, CV74, GI79, Mor70, Vui73]. simplified [GJS74]. simulation [Kos79c]. simulations [LS77]. simultaneity [MY78]. simultaneously [Coo79]. sixth [ACM74]. size [Con70, FMP75, MB72, Ruz79, SS78a, Sip79]. solution [Kat78, vL74]. solutions [PF77]. solvability [IK76]. Some [DL79, GJS74, GGJ76, Gre74, LJ71, Mun71, PS76, Yao79, DL74, IK76, Kos79a, Niv70, Rou70]. Sorting [TK76, Hor74, Fre75]. Space [PTC76, BS76, BM78, CLM76, Coo79, ET75, GLT79, Hor74, LT79, Pet79, Tom78]. specific [BC74]. specification [TWW78]. spectra [JS72]. speed [CH71]. speed-up [CH71]. speeding [AU72]. squared [Coo79]. stability [Mil74b]. stable [Hor74]. stack [PS77]. stacks [Pra73]. state [Raj71]. steps [Aus70, FMR79, KF77]. stochastic [San71]. Storage [CS74, RWG79, Coo73, Ros74]. store [Sud76]. Straight [Lyn78, AU70, Set72]. Straight-line [Lyn78]. strategy [Rei70]. String [Wag75, Gal76]. string-matching [Gal76]. String-to-String [Wag75]. strings [Ang79, KMR72]. Structure [Rob74, Ber71, ER75a, Fre79, Mey71, San71]. structured [Kos73]. structures [LED75, MS79, OC75, Pra75, RWG79]. structuring [Kim76, Kos76]. study [GJ73]. studying [MY78]. subdivision [LP76]. Subgraph [Lew78, HS76, LR78]. subject [Ano77]. Subrecursive [CM72b, Con70, Koz78, Meh74]. subset [Ter71]. substitution [Lew70a]. Subtree [O’D77]. super [LS76]. super-preconditioning [LS76]. sweeping [Sip79]. Switching [Weg78, TU79]. Symposium [ACM70, ACM71, ACM72, ACM73, ACM74, ACM75, ACM76, ACM77, ACM78, ACM79]. synchronization [Lip74, MY78]. synchronous [Gol78]. syntax [MV70]. syntax-directed [MV70]. system [HPS77, Kim76, Owi76, PF77, Wan76]. systems [Coo70, ERS78, Har77, Lad79, LR72, LB77, O’D77, Ros70a, ST77, Sha79, vL74].

table [Ter71]. Tape [BGW70, Hun73, LS77, Sud77]. Tape- [BGW70]. technique [AU72, Ber79, Fre75, GM77, Mil78]. techniques [TWW78]. tenth [ACM78]. test [BL75]. testers [CFG^{+}78]. Testing [Tar73, Tar74]. tests [Mil75, Sch72]. Texas [ACM73]. their [BM78, Tom78]. theorem [Boa71, Coo71, KF70, O’D79, Rei70]. theorem-proving [Coo71]. theorems [Bea77, Per75, Rou70, Ull70, Ros70a]. theoretic [GI78]. theories [Rac74, Rob74]. Theory [ACM70, ACM71, ACM72, ACM73, ACM74, ACM75, ACM76, ACM77, ACM78, ACM79, BM78, Con77, DL79, Man70, Mey71, O’D77, San71, ST78, Wag71]. thinking [Pra74]. third [ACM71]. Three [Ull70]. Time [CR72, Pet79, Tom78, BFP^{+}72, BGW70, BM78, Bur70, Coo72, Coo73, CS74, Coo79, Gal76, GHR76, Gre73, HW74, Hun73, JL74, Kos79c, Lad73, LLS74, LT79, LS77, Lon79, Pau77, SM73, Tar77, Tho79, Wot75]. Time-bounded [CR72, BGW70]. Time-space [Pet79, Tom78, BM78, LT79]. time-storage [Coo73]. toroidal [Fil78]. tours [WSB73]. tracing [Pag73]. trade [Coo73, Pet79]. trade-offs [Pet79]. tradeoffs [LT79, Tom78]. transducer [Per75]. transducers [ERS78, MV70, OR72, Raj71]. transduction [MV70]. transductions [Bak73]. transform [Fid72]. transformation [Ham74]. Transformations [AU70]. transition [Mor70]. transitive [Hir76]. translatability [Cha74]. Translating [Str70]. translations [CH71, LRS73, MY76]. traveling [WSB73, PS76]. Tree [Bak73, ERS78, Ros70a, Rou70, Ruz79, LJ71, MV70, Mey71, OR72, Per75, RWG79, Ell70]. tree-like [RWG79]. Tree-manipulating [Ros70a]. Tree-oriented [Rou70]. Tree-size [Ruz79]. trees [AHU73, AJ75, Fre75, KMR72, NR72]. trie [Bur76]. Turing [BGW70, Gil74, JS72]. turn [Val74]. Two [Fre75, Cha70, Con73, ERS78, Raj71, SS78a, Sch76]. two-dimensional [Cha70]. two-person [Sch76]. two-way [ERS78, Raj71]. Type [Con73, Fre79, TWW78].

undecidable [Cud70, CM72b]. unification [PW76]. unified [Gol78]. Uniformly [CGGG72]. unifying [O’D77]. units [LS77]. Universal [CW77, Rei79, Val76]. unrestricted [But71]. Unsolvability [Lew70b]. unusual [Wan73]. Upper [LT79]. upward [SG77]. useful [IK76]. using [Cha70, Tom78].

Validating [Set72]. validity [Gal75]. variables [Lip74]. vector [PRS74, ST77, vL74]. vector-addition [vL74]. verification [Mil73, Owi76]. Version [AU70, BC74, CR74, Coo75, Dow78, Kos76, Kos79a, Kos79b, Kos79c, LR70, Pip76, Rob71, ST77, Sim77, WS72]. versus [Lon79]. vertex [RT75]. via [Fid72]. VLSI [Tho79].

W [Gre74]. W-grammars [Gre74]. Washington [ACM74]. way [ERS78, Raj71, SS78a]. weak [Rac74, Rob74]. weights [Lue78]. whether [Fil78]. which [ET75, O’D79]. whose [Weg78]. without [ER75a]. Word [SM73]. Worst [GGU72]. Worst-case [GGU72].

X [Fre75].

References

ACM:1970:CRS

[ACM70] ACM, editor. Conference record of second annual ACM Symposium on Theory of Computing: papers presented at the symposium, Northampton, Massachusetts, May 4, 5, 6, 1970. ACM Press, New York, NY, USA, 1970. LCCN QA75.5 .A22 1970.

ACM:1971:CRT

[ACM71] ACM, editor. Conference record of third annual ACM Symposium on Theory of Computing: papers presented at the symposium, Shaker Heights, Ohio, May 3, 4, 5, 1971. ACM Press, New York, NY, USA, 1971. LCCN QA75.5 .A22 1971.

ACM:1972:CRF

[ACM72] ACM, editor. Conference record, Fourth Annual ACM Symposium on Theory of Computing: papers presented at the symposium, Denver, Colorado, May 1, 2, 3, 1972. ACM Press, New York, NY, USA, 1972. LCCN QA76.6 .A13 1972.

ACM:1973:CRF

[ACM73] ACM, editor. Conference record of Fifth Annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Austin, Texas, April 30–May 2, 1973. ACM Press, New York, NY, USA, 1973. LCCN QA76.6 .A16 1973.

ACM:1974:CRS

[ACM74] ACM, editor. Conference record of sixth annual ACM Symposium on Theory of Computing: papers presented at the symposium, Seattle, Washington, April 30–May 2, 1974. ACM Press, New York, NY, USA, 1974. LCCN QA76.6 .A13 1974.

ACM:1975:CRS

[ACM75] ACM, editor. Conference record of Seventh Annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Albuquerque, New Mexico, May 5–May 7, 1975. ACM Press, New York, NY, USA, 1975. LCCN QA76.6 .A16 1975.

ACM:1976:CRE

[ACM76] ACM, editor. Conference record of the eighth annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Hershey, Pennsylvania, May 3–5, 1976. ACM Press, New York, NY, USA, 1976. LCCN QA 76.6 A12 1976.

ACM:1977:CRN

[ACM77] ACM, editor. Conference record of the ninth annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Boulder, Colorado, May 2–4, 1977. ACM Press, New York, NY, USA, 1977. LCCN QA76.6 .A13 1977.

ACM:1978:CRT

[ACM78] ACM, editor. Conference record of the tenth annual ACM Symposium on Theory of Computing: papers presented at the Symposium, San Diego, California, May 1–3, 1978. ACM Press, New York, NY, USA, 1978. LCCN QA76.6.A13 1978.

ACM:1979:CRE

[ACM79] ACM, editor. Conference record of the eleventh annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Atlanta, Georgia, April 30–May 2, 1979. ACM Press, New York, NY, USA, 1979. LCCN QA75.5.A14 1979.

Aho:1973:FLC

[AHU73] A. V. Aho, J. E. Hopcroft, and J. D. Ullman. On finding lowest common ancestors in trees. In ACM [ACM73], pages 253–265. LCCN QA76.6 .A16 1973.

Aho:1975:OCG

[AJ75] A. V. Aho and S. C. Johnson. Optimal code generation for expression trees. In ACM [ACM75], pages 207– 217. LCCN QA76.6 .A16 1975.

Adleman:1977:RRI

[AM77] Leonard Adleman and Kenneth Manders. Reducibility, randomness, and intractibility (abstract). In ACM [ACM77], pages 151–163. LCCN QA76.6 .A13 1977.

Angluin:1979:FPC

[Ang79] Dana Angluin. Finding patterns common to a set of strings (extended abstract). In ACM [ACM79], pages 130–141. LCCN QA75.5.A14 1979.

Anonymous:1977:ASI

[Ano77] Anonymous. Author and subject index. In ACM [ACM77], pages 312– 314. LCCN QA76.6 .A13 1977.

Anonymous:1978:AFA

[Ano78] Anonymous. An analysis of the full alpha-beta pruning algorithm. In ACM [ACM78], pages 296–313. LCCN QA76.6.A13 1978.

Aho:1970:TSL

[AU70] A. V. Aho and J. D. Ullman.

Transformations on straight line programs — (preliminary version). In ACM [ACM70], pages 136–148. LCCN QA75.5 .A22 1970.

Aho:1971:CFL

[AU71] Alfred V. Aho and Jeffrey D. Ullman. The care and feeding of LR(k) grammars. In ACM [ACM71], pages 159–170. LCCN QA75.5 .A22 1971.

Aho:1972:TSL

[AU72] Alfred V. Aho and Jeffrey D. Ullman. A technique for speeding up LR(k) parsers. In ACM [ACM72], pages 251–263. LCCN QA76.6 .A13 1972.

Aho:1975:NLR

[AU75] A. V. Aho and J. D. Ullman. Node listings for reducible flow graphs. In ACM [ACM75], pages 177–185. LCCN QA76.6 .A16 1975.

Ausiello:1970:BNS

[Aus70] Giorgio Ausiello. On bounds on the number of steps to compute functions. In ACM [ACM70], pages 41– 47. LCCN QA75.5 .A22 1970.

Angluin:1977:FPA

[AV77] Dana Angluin and Leslie G. Valiant.

Fast probabilistic algorithms for Hamiltonian circuits and matchings. In ACM [ACM77], pages 30–41. LCCN QA76.6 .A13 1977.

Baker:1973:TTF

[Bak73] Brenda S. Baker. Tree transductions and families of tree languages.

In ACM [ACM73], pages 200–206. LCCN QA76.6 .A16 1973.

Borodin:1974:NAC

[BC74] Allan Borodin and Stephen Cook. On the number of additions to compute specific polynomials (preliminary version). In ACM [ACM74], pages 342–347. LCCN QA76.6 .A13 1974.

Brockett:1973:OES

[BD73] Roger W. Brockett and David Dobkin. On the optimal evaluation of a set of bilinear forms. In ACM [ACM73], pages 88–95. LCCN QA76.6 .A16 1973.

Beatty:1977:ITL

[Bea77] John C. Beatty. Iteration theorems for LL(k) languages (extended abstract). In ACM [ACM77], pages 122–131. LCCN QA76.6 .A13 1977.

Berry:1971:BSE

[Ber71] Daniel M. Berry. Block structure (extended abstract): Retention or deletion? In ACM [ACM71], pages 86–100. LCCN QA75.5 .A22 1971.

Berman:1979:CTA

[Ber79] Francine Berman. A completeness technique for d-axiomatizable semantics. In ACM [ACM79], pages 160–166. LCCN QA75.5.A14 1979.

Blum:1972:LTB

[BFP^{+}72] Manuel Blum, Robert W. Floyd, Vaughan Pratt, Ronald L. Rivest, and Robert E. Tarjan. Linear time bounds for median computations. In ACM [ACM72], pages 119–124. LCCN QA76.6 .A13 1972.

Book:1970:TTB

[BGW70] Ronald V. Book, Sheila A.

Greibach, and Ben Wegbreit. Tapeand time-bounded Turing acceptors and AFLs (extended abstract). In ACM [ACM70], pages 92–99. LCCN QA75.5 .A22 1970.

Booth:1975:LAR

[BL75] Kellogg S. Booth and George S.

Lueker. Linear algorithms to recognize interval graphs and test for the consecutive ones property. In ACM [ACM75], pages 255–265. LCCN QA76.6 .A16 1975.

Bruss:1978:TSC

[BM78] Anni R. Bruss and Albert R. Meyer.

On time-space classes and their relation to the theory of real addition. In ACM [ACM78], pages 233–239. LCCN QA76.6.A13 1978.

Beeri:1979:ERD

[BMSU79] Catriel Beeri, Alberto O. Mendelzon, Yehoshua Sagiv, and Jeffrey D. Ullman. Equivalence of relational database schemes. In ACM [ACM79], pages 319–329. LCCN QA75.5.A14 1979.

Book:1974:ILC

[BNP74] Ronald Book, Maurice Nivat, and Michael Paterson. Intersections of linear context-free languages and reversal-bounded multipushdown machines (extended abstract). In ACM [ACM74], pages 290–296. LCCN QA76.6 .A13 1974.

Boasson:1971:ITO

[Boa71] L. Boasson. An iteration theorem for one-counter languages. In ACM [ACM71], pages 116–120. LCCN QA75.5 .A22 1971.

Brown:1977:CPQ

[Bro77] Mark R. Brown. The complexity of priority queue maintenance. In ACM [ACM77], pages 42–48. LCCN QA76.6 .A13 1977.

Bentley:1976:DCM

[BS76] Jon Louis Bentley and Michael Ian Shamos. Divide-and-conquer in multidimensional space. In ACM [ACM76], pages 220–230. LCCN QA 76.6 A12 1976.

Brown:1978:RLL

[BT78] Mark R. Brown and Robert E. Tarjan. A representation for linear lists with movable fingers. In ACM [ACM78], pages 19–29. LCCN QA76.6.A13 1978.

Burkhard:1970:CPR

[Bur70] W. A. Burkhard. Complexity problems in real time computation. In ACM [ACM70], pages 62–69. LCCN QA75.5 .A22 1970.

Burkhard:1976:ART

[Bur76] Walter A. Burkhard. Associative retrieval trie hash-coding (extended abstract). In ACM [ACM76], pages 211–219. LCCN QA 76.6 A12 1976.

Buttelmann:1971:GFA

[But71] H. W. Buttelmann. On generalized finite automata and unrestricted generative grammars. In ACM [ACM71], pages 63–77. LCCN QA75.5 .A22 1971.

Bass:1970:HBC

[BY70] L. J. Bass and Paul R. Young. Hierarchies based on computational complexity and irregularities of class determining measured sets (preliminary report). In ACM [ACM70], pages 37–40. LCCN QA75.5 .A22 1970.

Chazelle:1979:DPC

[CD79] Bernard Chazelle and David Dobkin.

Decomposing a polygon into its convex parts. In ACM [ACM79], pages 38–48. LCCN QA75.5.A14 1979.

Carter:1978:EAM

[CFG^{+}78] Larry Carter, Robert Floyd, John Gill, George Markowsky, and Mark Wegman. Exact and approximate membership testers. In ACM [ACM78], pages 59–65. LCCN QA76.6.A13 1978.

Carlyle-Greibach:1972:UEA

[CGGG72] Sheila Carlyle-Greibach, Seymour Ginsburg, and Jonathan Goldstine. Uniformly erasable AFL. In ACM [ACM72], pages 207–213. LCCN QA76.6 .A13 1972.

Constable:1971:CFT

[CH71] R. L. Constable and J. Hartmanis. Complexity of formal translations and speed-up results. In ACM [ACM71], pages 244–250. LCCN QA75.5 .A22 1971.

Chandra:1979:CQR

[CH79] Ashok K. Chandra and David Harel.

Computable queries for relational data bases (preliminary report). In ACM [ACM79], pages 309–318. LCCN QA75.5.A14 1979.

Chang:1970:ATD

[Cha70] Shi-Kuo Chang. The analysis of two-dimensional patterns using picture processing grammars. In ACM [ACM70], pages 206–216. LCCN QA75.5 .A22 1970.

Chang:1971:PCL

[Cha71] Shi-Kuo Chang. On the parallel computation of local operations. In ACM [ACM71], pages 101–115. LCCN QA75.5 .A22 1971.

Chandra:1974:DTC

[Cha74] Ashok K. Chandra. Degrees of translatability and canonical forms in program schemas: Part I. In ACM [ACM74], pages 1–12. LCCN QA76.6 .A13 1974.

Cardoza:1976:ESC

[CLM76] E. Cardoza, R. Lipton, and A. R. Meyer. Exponential space complete problems for Petri nets and commutative semigroups (preliminary report). In ACM [ACM76], pages 50– 54. LCCN QA 76.6 A12 1976.

Chandra:1972:PSE

[CM72a] Ashok K. Chandra and Zohar Manna. Program schemas with equality. In ACM [ACM72], pages 52–64. LCCN QA76.6 .A13 1972.

Constable:1972:SPS

[CM72b] Robert L. Constable and Steven S. Muchnick. Subrecursive program schemata I & II (I. Undecidable equivalence problems, II. Decidable equivalence problems). In ACM [ACM72], pages 1–17. LCCN QA76.6 .A13 1972.

Chandra:1977:OIC

[CM77] Ashok K. Chandra and Philip M.

Merlin. Optimal implementation of conjunctive queries in relational data bases. In ACM [ACM77], pages 77–90. LCCN QA76.6 .A13 1977.

Constable:1970:SPS

[Con70] Robert L. Constable. On the size of programs in subrecursive formalisms. In ACM [ACM70], pages 1–9. LCCN QA75.5 .A22 1970.

Constable:1971:LS

[Con71] Robert L. Constable. Loop schemata. In ACM [ACM71], pages 24–39. LCCN QA75.5 .A22 1971.

Constable:1973:TTC

[Con73] Robert L. Constable. Type two computational complexity. In ACM [ACM73], pages 108–121. LCCN QA76.6 .A16 1973.

Constable:1977:TPL

[Con77] Robert L. Constable. On the theory of programming logics. In ACM [ACM77], pages 269–285. LCCN QA76.6 .A13 1977.

Cook:1970:PSL

[Coo70] Stephen A. Cook. Path systems and language recognition. In ACM [ACM70], pages 70–72. LCCN QA75.5 .A22 1970.

Cook:1971:CTP

[Coo71] Stephen A. Cook. The complexity of theorem-proving procedures. In ACM [ACM71], pages 151–158. LCCN QA75.5 .A22 1971.

Cook:1972:HNT

[Coo72] Stephen A. Cook. A hierarchy for nondeterministic time complexity. In ACM [ACM72], pages 187–192. LCCN QA76.6 .A13 1972.

Cook:1973:OTS

[Coo73] Stephen A. Cook. An observation on time-storage trade off. In ACM [ACM73], pages 29–33. LCCN QA76.6 .A16 1973.

Cook:1975:FCP

[Coo75] Stephen A. Cook. Feasibly constructive proofs and the propositional calculus (preliminary version). In ACM [ACM75], pages 83–97. LCCN QA76.6 .A16 1975.

Cook:1979:DCA

[Coo79] Stephen A. Cook. Deterministic CFL’s are accepted simultaneously in polynomial time and log squared space. In ACM [ACM79], pages 338–345. LCCN QA75.5.A14 1979.

Cook:1972:TBR

[CR72] Stephen A. Cook and Robert A. Reckhow. Time-bounded random access machines. In ACM [ACM72], pages 73–80. LCCN QA76.6 .A13 1972.

Cook:1974:LPP

[CR74] Stephen Cook and Robert Reckhow. On the lengths of proofs in the propositional calculus (preliminary version). In ACM [ACM74], pages 135–148. LCCN QA76.6 .A13 1974.

Cook:1974:SRD

[CS74] Stephen Cook and Ravi Sethi. Storage requirements for deterministic/ polynomial time recognizable languages. In ACM [ACM74], pages 33–39. LCCN QA76.6 .A13 1974.

Case:1978:AHM

[CS78] John Case and Carl Smith.

Anomaly hierarchies of mechanized inductive inference. In ACM [ACM78], pages 314–319. LCCN QA76.6.A13 1978.

Cudia:1970:DHU

[Cud70] Dennis F. Cudia. The degree hierarchy of undecidable problems of formal grammars. In ACM [ACM70], pages 10–21. LCCN QA75.5 .A22 1970.

Courcelle:1974:SAS

[CV74] B. Courcelle and J. Vuillemin. Semantics and axiomatics of a simple recursive language. In ACM [ACM74], pages 13–26. LCCN QA76.6 .A13 1974.

Carter:1977:UCH

[CW77] J. Lawrence Carter and Mark N.

Wegman. Universal classes of hash functions (extended abstract). In ACM [ACM77], pages 106–112. LCCN QA76.6 .A13 1977.

Yao:1976:ABS

[cY76] Andrew Chi chih Yao. On the average behavior of set merging algorithms (extended abstract). In ACM [ACM76], pages 192–195. LCCN QA 76.6 A12 1976.

Dobkin:1974:SGB

[DL74] David Dobkin and R. J. Lipton. On some generalizations of binary search. In ACM [ACM74], pages 310–316. LCCN QA76.6 .A13 1974.

DeMillo:1979:SCB

[DL79] Richard A. DeMillo and Richard J. Lipton. Some connections between mathematical logic and complexity theory. In ACM [ACM79], pages 153–159. LCCN QA75.5.A14 1979.

Doeppner:1976:APP

[Doe76] Thomas W. Doeppner. On abstractions of parallel programs. In ACM [ACM76], pages 65–72. LCCN QA 76.6 A12 1976.

Dowd:1978:PRA

[Dow78] Martin Dowd. Propositional representation of arithmetic proofs (preliminary version). In ACM [ACM78], pages 246–252. LCCN QA76.6.A13 1978.

Egli:1975:CCP

[EC75] Herbert Egli and Robert L. Constable. Computability concepts for programming language semantics. In ACM [ACM75], pages 98–106. LCCN QA76.6 .A16 1975.

Eilenberg:1973:CSC

[Eil73] Samuel Eilenberg. Classes of semigroups and classes of sets. In ACM [ACM73], pages 266–267. LCCN QA76.6 .A16 1973.

Ellis:1970:PTA

[Ell70] Clarence A. Ellis. Probabilistic Tree Automata. In ACM [ACM70], pages 198–205. LCCN QA75.5 .A22 1970.

Ehrenfeucht:1975:CSL

[ER75a] A. Ehrenfeucht and G. Rozenberg.

On the (combinatorial) structure of L languages without interactions (extended abstract). In ACM [ACM75], pages 137–144. LCCN QA76.6 .A16 1975.

Ehrenfeucht:1975:PFL

[ER75b] A. Ehrenfeucht and G. Rozenberg. On (un)predictability of formal languages (extended abstract). In ACM [ACM75], pages 117–120. LCCN QA76.6 .A16 1975.

Engelfriet:1978:TTS

[ERS78] J. Engelfriet, G. Rozenberg, and G. Slutzki. Tree transducers, L systems and two-way machines (extended abstract). In ACM [ACM78], pages 66–74. LCCN QA76.6.A13 1978.

Even:1975:CPW

[ET75] Shimon Even and R. Endre Tarjan.

A combinatorial problem which is complete in polynomial space. In ACM [ACM75], pages 66–71. LCCN QA76.6 .A16 1975.

Ehrenfeucht:1974:CMR

[EZ74] Andrzej Ehrenfeucht and Paul Zeiger. Complexity measures for regular expressions. In ACM [ACM74], pages 75–79. LCCN QA76.6 .A13 1974.

Flajolet:1979:CIC

[FFV79] P. Flajolet, J. Françon, and J. Vuillemin. Computing integrated costs of sequences of operations with application to dictionaries. In ACM [ACM79], pages 49–61. LCCN QA75.5.A14 1979.

Fiduccia:1971:FMM

[Fid71] Charles M. Fiduccia. Fast matrix multiplication. In ACM [ACM71], pages 45–49. LCCN QA75.5 .A22 1971.

Fiduccia:1972:PED

[Fid72] Charles M. Fiduccia. Polynomial evaluation via the division algorithm the fast Fourier transform revisited. In ACM [ACM72], pages 88–93. LCCN QA76.6 .A13 1972.

Filotti:1978:EAD

[Fil78] I. S. Filotti. An efficient algorithm for determining whether a cubic graph is toroidal. In ACM [ACM78], pages 133–142. LCCN QA76.6.A13 1978.

Fischer:1977:PML

[FL77] Michael J. Fischer and Richard E. Ladner. Propositional modal logic of programs. In ACM [ACM77], pages 286–294. LCCN QA76.6 .A13 1977.

Fischer:1975:LBS

[FMP75] Michael J. Fischer, Albert R. Meyer, and Michael S. Paterson. Lower bounds on the size of Boolean formulas: Preliminary Report. In ACM [ACM75], pages 37–44. LCCN QA76.6 .A16 1975.

Filotti:1979:DGG

[FMR79] I. S. Filotti, Gary L. Miller, and John Reif. On determining the genus of a graph in O(v^{O(g)}) steps (preliminary report). In ACM [ACM79], pages 27–37. LCCN QA75.5.A14 1979.

Fredman:1975:TAP

[Fre75] Michael L. Fredman. Two applications of a probabilistic search technique: Sorting X+Y and building balanced search trees. In ACM [ACM75], pages 240–244. LCCN QA76.6 .A16 1975.

Fredman:1979:NOD

[Fre79] Michael L. Fredman. A near optimal data structure for a type of range query problem. In ACM [ACM79], pages 62–66. LCCN QA75.5.A14 1979.

Fischer:1973:FLI

[FS73] Michael J. Fischer and Larry J.

Stockmeyer. Fast on-line integer multiplication. In ACM [ACM73], pages 67–72. LCCN QA76.6 .A16 1973.

Fong:1976:FDF

[FU76] Amelia C. Fong and Jeffrey D. Ullman. Finding the depth of a flow graph. In ACM [ACM76], pages 121–125. LCCN QA 76.6 A12 1976.

Fortune:1978:PRA

[FW78] Steven Fortune and James Wyllie.

Parallelism in random access machines. In ACM [ACM78], pages 114–118. LCCN QA76.6.A13 1978.

Galil:1975:VCB

[Gal75] Zvi Galil. On the validity and complexity of bounded resolution. In ACM [ACM75], pages 72–82. LCCN QA76.6 .A16 1975.

Galil:1976:RTA

[Gal76] Zvi Galil. Real-time algorithms for string-matching and palindrome recognition. In ACM [ACM76], pages 161–173. LCCN QA 76.6 A12 1976.

Ginsburg:1971:ICF

[GG71] Seymour Ginsburg and Jonathan Goldstine. Intersection-closed full AFL and the recursively enumerable languages. In ACM [ACM71], pages 121–131. LCCN QA75.5 .A22 1971.

Garey:1976:SNC

[GGJ76] M. R. Garey, R. L. Graham, and D. S. Johnson. Some NP-complete geometric problems. In ACM [ACM76], pages 10–22. LCCN QA 76.6 A12 1976.

Garey:1972:WCA

[GGU72] M. R. Garey, R. L. Graham, and J. D. Ullman. Worst-case analysis of memory allocation algorithms. In ACM [ACM72], pages 143–150. LCCN QA76.6 .A13 1972.

Graham:1976:LCF

[GHR76] Susan L. Graham, Michael A. Harrison, and Walter L. Ruzzo. On line context free language recognition in less than cubic time (extended abstract). In ACM [ACM76], pages 112–120. LCCN QA 76.6 A12 1976.

Gurari:1978:NCN

[GI78] Eitan M. Gurari and Oscar H. Ibarra. An NP-complete numbertheoretic problem. In ACM [ACM78], pages 205–215. LCCN QA76.6.A13 1978.

Gurari:1979:CEP

[GI79] Eitan M. Gurari and Oscar H.

Ibarra. The complexity of the equivalence problem for counter machines, semilinear sets, and simple programs. In ACM [ACM79], pages 142–152. LCCN QA75.5.A14 1979.

Gill:1974:CCP

[Gil74] John T. Gill. Computational complexity of probabilistic Turing machines. In ACM [ACM74], pages 91– 95. LCCN QA76.6 .A13 1974.

Gentleman:1973:AAC

[GJ73] W. M. Gentleman and S. C. Johnson. Analysis of algorithms, a case study: Determinants of polynomials. In ACM [ACM73], pages 135– 141. LCCN QA76.6 .A16 1973.

Garey:1974:SSN

[GJS74] M. R. Garey, D. S. Johnson, and L. Stockmeyer. Some simplified NP-complete problems. In ACM [ACM74], pages 47–63. LCCN QA76.6 .A13 1974.

Gabow:1978:AEC

[GK78] Harold N. Gabow and Oded Kariv.

Algorithms for edge coloring bipartite graphs. In ACM [ACM78], pages 184–192. LCCN QA76.6.A13 1978.

Garland:1972:ES

[GL72] Stephen J. Garland and David C.

Luckham. On the equivalence of schemes. In ACM [ACM72], pages 65–72. LCCN QA76.6 .A13 1972.

Ginsburg:1975:CCG

[GL75] Seymour Ginsburg and Nancy Lynch. Comparative complexity of grammar forms. In ACM [ACM75], pages 153–158. LCCN QA76.6 .A16 1975.

Gilbert:1979:PPC

[GLT79] John R. Gilbert, Thomas Lengauer, and Robert Endre Tarjan. The pebbling problem is complete in polynomial space. In ACM [ACM79], pages 237–248. LCCN QA75.5.A14 1979.

Gonnet:1977:AIH

[GM77] Gaston Gonnet and Ian Munro. The analysis of an improved hashing technique. In ACM [ACM77], pages 113–121. LCCN QA76.6 .A13 1977.

Guibas:1977:NRL

[GMPR77] Leo J. Guibas, Edward M. Mc-Creight, Michael F. Plass, and Janet R. Roberts. A new representation for linear lists. In ACM [ACM77], pages 49–60. LCCN QA76.6 .A13 1977.

Galil:1979:NFG

[GN79] Zvi Galil and Amnon Naamad. Network flow and generalized path compression. In ACM [ACM79], pages 13–26. LCCN QA75.5.A14 1979.

Goldschlager:1978:UAM

[Gol78] Leslie M. Goldschlager. A unified approach to models of synchronous parallel machines. In ACM [ACM78], pages 89–94. LCCN QA76.6.A13 1978.

Greibach:1973:JPD

[Gre73] Sheila A. Greibach. Jump PDA’s, deterministic context-free languages principal AFDLs and polynomial time recognition — (extended abstract). In ACM [ACM73], pages 20–28. LCCN QA76.6 .A16 1973.

Greibach:1974:SRW

[Gre74] Sheila A. Greibach. Some restrictions on W-grammars. In ACM [ACM74], pages 256–265. LCCN QA76.6 .A13 1974.

Guibas:1976:ADH

[GS76] Leo J. Guibas and Endre Szemeredi.

The analysis of double hashing (extended abstract). In ACM [ACM76], pages 187–191. LCCN QA 76.6 A12 1976.

Hennessy:1977:PPM

[HA77] M. C. B. Hennessy and E. A.

Ashcroft. Parameter-passing mechanisms and nondeterminism. In ACM [ACM77], pages 306–311. LCCN QA76.6 .A13 1977.

Hamlet:1972:PPA

[Ham72] R. G. Hamlet. A patent problem for abstract programming languages; machine-independent computations. In ACM [ACM72], pages 193–197. LCCN QA76.6 .A13 1972.

Hammer:1974:NGT

[Ham74] Michael Hammer. A new grammatical transformation into LL(k) form (extended abstract). In ACM [ACM74], pages 266–275. LCCN QA76.6 .A13 1974.

Hartmanis:1977:RBD

[Har77] J. Hartmanis. Relations between diagonalization, proof systems, and complexity gaps. In ACM [ACM77], pages 223–227. LCCN QA76.6 .A13 1977.

Hartmanis:1976:IDN

[HB76] J. Hartmanis and L. Berman. On isomorphisms and density of NP and other complete sets. In ACM [ACM76], pages 30–40. LCCN QA 76.6 A12 1976.

Hirschberg:1976:PAT

[Hir76] D. S. Hirschberg. Parallel algorithms for the transitive closure and the connected component problems. In ACM [ACM76], pages 55–57. LCCN QA 76.6 A12 1976.

Hyafil:1975:CPE

[HK75] L. Hyafil and H. T. Kung. The complexity of parallel evaluation of linear recurrence. In ACM [ACM75], pages 12–22. LCCN QA76.6 .A16 1975.

Hopcroft:1973:DAC

[HM73] J. Hopcroft and J. Musinski. Duality applied to the complexity of matrix multiplications and other bilinear forms. In ACM [ACM73], pages 73–87. LCCN QA76.6 .A16 1973.

Harel:1977:CCL

[HMP77] D. Harel, A. R. Meyer, and V. R. Pratt. Computability and completeness in logics of programs (preliminary report). In ACM [ACM77], pages 261–268. LCCN QA76.6 .A13 1977.

Horowitz:1972:ARF

[Hor72] Ellis Horowitz. Algorithms for rational function arithmetic operations. In ACM [ACM72], pages 108–118. LCCN QA76.6 .A13 1972.

Horvath:1974:ESS

[Hor74] Edward C. Horvath. Efficient stable sorting with minimal extra space. In ACM [ACM74], pages 194–215. LCCN QA76.6 .A13 1974.

Harel:1977:CAS

[HPS77] David Harel, Amir Puneli, and Jonathan Stavi. A complete axiomatic system for proving deductions about recursive programs. In ACM [ACM77], pages 249–260. LCCN QA76.6 .A13 1977.

Hyafil:1974:EAC

[HPV74] L. Hyafil, F. Prusker, and J. Vuillemin.

An efficient algorithm for computing optimal disk merge patterns. (extended abstract). In ACM [ACM74], pages 216–229. LCCN QA76.6 .A13 1974.

Hunt:1974:CPB

[HR74] H. B. Hunt and D. J. Rosenkrantz.

Computational parallels between the regular and context-free languages. In ACM [ACM74], pages 64–74. LCCN QA76.6 .A13 1974.

Hunt:1975:CGR

[HS75] H. B. Hunt and T. G. Szymanski. On the complexity of grammar and related problems. In ACM [ACM75], pages 54–65. LCCN QA76.6 .A16 1975.

Hunt:1976:DRF

[HS76] H. B. Hunt and T. G. Szymanski. Dichotomization, reachability, and the forbidden subgraph problem (extended abstract). In ACM [ACM76], pages 126–134. LCCN QA 76.6 A12 1976.

Hecht:1972:FGR

[HU72] Matthew S. Hecht and Jeffrey D. Ullman. Flow graph reducibility. In ACM [ACM72], pages 238–250. LCCN QA76.6 .A13 1972.

Hunt:1973:TTC

[Hun73] H. B. Hunt. On the time and tape complexity of languages I. In ACM [ACM73], pages 10–19. LCCN QA76.6 .A16 1973.

Hopcroft:1974:LTA

[HW74] J. E. Hopcroft and J. K. Wong. Linear time algorithm for isomorphism of planar graphs (preliminary report). In ACM [ACM74], pages 172– 184. LCCN QA76.6 .A13 1974.

Hyafil:1978:PEM

[Hya78] Laurent Hyafil. On the parallel evaluation of multivariate polynomials. In ACM [ACM78], pages 193–195. LCCN QA76.6.A13 1978.

Ibarra:1976:UDS

[IK76] Oscar H. Ibarra and Chul E. Kim. A useful device for showing the solvability of some decision problems. In ACM [ACM76], pages 135–140. LCCN QA 76.6 A12 1976.

Itai:1977:FMC

[IR77] Alon Itai and Michael Rodeh. Finding a minimum circuit in a graph. In ACM [ACM77], pages 1–10. LCCN QA76.6 .A13 1977.

Ja:1978:OEP

[Ja’78] Joseph Ja’ Ja’. Optimal evaluation of pairs of bilinear forms. In ACM [ACM78], pages 173–183. LCCN QA76.6.A13 1978.

JaJa:1979:CBF

[Ja’79] Joseph Ja’Ja’. On the complexity of bilinear forms with commutativity. In ACM [ACM79], pages 197–208. LCCN QA75.5.A14 1979.

Jones:1974:CPD

[JL74] Neil D. Jones and William T.

Laaser. Complete problems for deterministic polynomial time. In ACM [ACM74], pages 40–46. LCCN QA76.6 .A13 1974.

Johnson:1973:AAC

[Joh73] David S. Johnson. Approximation algorithms for combinatorial problems. In ACM [ACM73], pages 38– 49. LCCN QA76.6 .A16 1973.

Jones:1972:TMS

[JS72] Neil D. Jones and Alan L. Selman.

Turing machines and the spectra of first-order formulas with equality. In ACM [ACM72], pages 157–167. LCCN QA76.6 .A13 1972.

Katseff:1978:NSC

[Kat78] Howard P. Katseff. A new solution to the critical section problem. In ACM [ACM78], pages 86–88. LCCN QA76.6.A13 1978.

Kedem:1974:CDR

[Ked74] Zvi M. Kedem. Combining dimensionality and rate of growth arguments for establishing lower bounds on the number of multiplications. In ACM [ACM74], pages 334–341. LCCN QA76.6 .A13 1974.

King:1970:IOT

[KF70] James C. King and Robert W. Floyd. An interpretation oriented theorem prover over integers. In ACM [ACM70], pages 169–179. LCCN QA75.5 .A22 1970.

Kintala:1977:CRN

[KF77] Chandra M. R. Kintala and Patrick C. Fischer. Computations with a restricted number of nondeterministic steps (extended abstract). In ACM [ACM77], pages 178–185. LCCN QA76.6 .A13 1977.

Kirkpatrick:1978:CGM

[KH78] David G. Kirkpatrick and Pavol Hell. On the completeness of a generalized matching problem. In ACM [ACM78], pages 240–245. LCCN QA76.6.A13 1978.

Kimura:1976:ASP

[Kim76] Takayuki Kimura. An algebraic system for process structuring and interprocess communication. In ACM [ACM76], pages 92–100. LCCN QA 76.6 A12 1976.

Kirkpatrick:1972:ANC

[Kir72] David Kirkpatrick. On the additions necessary to compute certain functions. In ACM [ACM72], pages 94– 101. LCCN QA76.6 .A13 1972.

Kirkpatrick:1974:DGP

[Kir74] David Kirkpatrick. Determining graph properties from matrix representations. In ACM [ACM74], pages 84–90. LCCN QA76.6 .A13 1974.

Karp:1972:RIR

[KMR72] Richard M. Karp, Raymond E.

Miller, and Arnold L. Rosenberg. Rapid identification of repeated patterns in strings, trees and arrays. In ACM [ACM72], pages 125–136. LCCN QA76.6 .A13 1972.

Kosaraju:1973:ASP

[Kos73] S. Rao Kosaraju. Analysis of structured programs. In ACM [ACM73], pages 240–252. LCCN QA76.6 .A16 1973.

Kosaraju:1976:SFP

[Kos76] S. Rao Kosaraju. On structuring flowcharts (preliminary version). In ACM [ACM76], pages 101–111. LCCN QA 76.6 A12 1976.

Kosaraju:1979:FPP

[Kos79a] S. Rao Kosaraju. Fast parallel processing array algorithms for some graph problems (preliminary version). In ACM [ACM79], pages 231– 236. LCCN QA75.5.A14 1979.

Kosaraju:1979:MSP

[Kos79b] S. Rao Kosaraju. On a multidimensional search problem (preliminary version). In ACM [ACM79], pages 67–73. LCCN QA75.5.A14 1979.

Kosaraju:1979:RTS

[Kos79c] S. Rao Kosaraju. Real-time simulation of concatenable double-ended queues by double-ended queues (preliminary version). In ACM [ACM79], pages 346–351. LCCN QA75.5.A14 1979.

Kozen:1977:CFP

[Koz77] Dexter Kozen. Complexity of finitely presented algebras. In ACM [ACM77], pages 164–177. LCCN QA76.6 .A13 1977.

Kozen:1978:ISC

[Koz78] Dexter Kozen. Indexing of subrecursive classes. In ACM [ACM78], pages 287–295. LCCN QA76.6.A13 1978.

Kung:1972:BME

[Kun72] H. T. Kung. A bound on the multiplication efficiency of iteration. In ACM [ACM72], pages 102–107. LCCN QA76.6 .A13 1972.

Kung:1973:CCA

[Kun73] H. T. Kung. The computational complexity of algebraic numbers. In ACM [ACM73], pages 152–159. LCCN QA76.6 .A16 1973.

Kung:1974:NAL

[Kun74] H. T. Kung. New algorithms and lower bounds for the parallel evaluation of certain rational expressions. In ACM [ACM74], pages 323–333. LCCN QA76.6 .A13 1974.

Ladner:1973:PTR

[Lad73] Richard E. Ladner. Polynominal time reducibility. In ACM [ACM73], pages 122–129. LCCN QA76.6 .A16

Ladner:1979:CPS

[Lad79] Richard E. Ladner. The complexity of problems in systems of communicating sequential processes (extended abstract). In ACM [ACM79], pages 214–223. LCCN QA75.5.A14 1979.

Lynch:1977:ERB

[LB77] Nancy A. Lynch and Edward K.

Blum. Efficient reducibility between programming systems (preliminary report). In ACM [ACM77], pages 228–238. LCCN QA76.6 .A13 1977.

Lipton:1975:CMH

[LD75] Richard J. Lipton and David Dobkin. Complexity measures and hierarchies for the evaluation of integers, polynomials, and n-linear forms. In ACM [ACM75], pages 1–5. LCCN QA76.6 .A16 1975.

Lipton:1975:CCS

[LED75] R. J. Lipton, S. C. Eisenstat, and R. A. DeMillo. The complexity of control structures and data structures. In ACM [ACM75], pages 186– 193. LCCN QA76.6 .A16 1975.

Lewis:1970:CFL

[Lew70a] David J. Lewis. Closure of families of languages under substitution operators. In ACM [ACM70], pages 100–108. LCCN QA75.5 .A22 1970.

Lewis:1970:UCC

[Lew70b] F. D. Lewis. Unsolvability considerations in computational complexity. In ACM [ACM70], pages 22–30. LCCN QA75.5 .A22 1970.

Lewis:1978:CMS

[Lew78] John M. Lewis. On the complexity of the Maximum Subgraph Problem. In ACM [ACM78], pages 265– 274. LCCN QA76.6.A13 1978.

Lindstrom:1970:DPI

[Lin70] G. Lindstrom. The design of parsers for incremental language processors. In ACM [ACM70], pages 81–91. LCCN QA75.5 .A22 1970.

Lipton:1974:LSP

[Lip74] R. J. Lipton. Limitations of synchronization primitives with conditional branching and global variables. In ACM [ACM74], pages 230– 241. LCCN QA76.6 .A13 1974.

Levy:1971:SRT

[LJ71] L. S. Levy and A. K. Joshi. Some results in tree automata. In ACM [ACM71], pages 78–85. LCCN QA75.5 .A22 1971.

Ladner:1974:CPT

[LLS74] Richard Ladner, Nancy Lynch, and Alan Selman. Comparison of polynomial-time reducibilities. In ACM [ACM74], pages 110–121. LCCN QA76.6 .A13 1974.

Lynch:1973:SDH

[LMF73] Nancy Lynch, Albert Meyer, and Michael Fischer. Sets that don’t help. In ACM [ACM73], pages 130– 134. LCCN QA76.6 .A16 1973.

Long:1979:GRV

[Lon79] Timothy J. Long. On γ-reducibility versus polynomial time many-one reducibility (extended abstract). In ACM [ACM79], pages 278–287. LCCN QA75.5.A14 1979.

Lee:1976:LPP

[LP76] D. T. Lee and F. P. Preparata. Location of a point in a planar subdivision and its applications. In ACM [ACM76], pages 231–235. LCCN QA 76.6 A12 1976.

Landweber:1970:RPA

[LR70] L. H. Landweber and E. L. Robertson. Recursive properties of abstract complexity classes (preliminary version). In ACM [ACM70], pages 31–36. LCCN QA75.5 .A22 1970.

Lindenmayer:1972:DSL

[LR72] A. Lindenmayer and G. Rozenberg. Developmental systems and languages. In ACM [ACM72], pages 214–221. LCCN QA76.6 .A13 1972.

LaPaugh:1978:SHP

[LR78] Andrea S. LaPaugh and Ronald L.

Rivest. The subgraph homeomorphism problem. In ACM [ACM78], pages 40–50. LCCN QA76.6.A13 1978.

Lewis:1973:ATE

[LRS73] P. M. Lewis, D. J. Rosenkrantz, and R. E. Stearns. Attributed translations (extended abstract). In ACM [ACM73], pages 160–171. LCCN QA76.6 .A16 1973.

Lipton:1976:EPS

[LS76] Richard J. Lipton and Larry J.

Stockmeyer. Evaluation of polynomials with super-preconditioning. In ACM [ACM76], pages 174–180. LCCN QA 76.6 A12 1976.

Leong:1977:NRT

[LS77] Benton Leong and Joel Seiferas. New real-time simulations of multihead tape units. In ACM [ACM77], pages 239–248. LCCN QA76.6 .A13 1977.

Lengauer:1979:ULB

[LT79] Thomas Lengauer and Robert Endre Tarjan. Upper and lower bounds on time-space tradeoffs. In ACM [ACM79], pages 262–277. LCCN QA75.5.A14 1979.

Lueker:1978:MPG

[Lue78] George S. Lueker. Maximization problems on graphs with edge weights chosen from a normal distribution (extended abstract). In ACM [ACM78], pages 13–18. LCCN QA76.6.A13 1978.

Lynch:1978:SLP

[Lyn78] Nancy A. Lynch. Straight-line program length as a parameter for complexity measures. In ACM [ACM78], pages 150–161. LCCN QA76.6.A13 1978.

Manders:1976:NCD

[MA76] Kenneth Manders and Leonard Adleman. NP-complete decision problems for quadratic polynomials. In ACM [ACM76], pages 23–29. LCCN QA 76.6 A12 1976.

Machtey:1971:CCF

[Mac71] Michael Machtey. Classification of computable functions by primitive recursive classes. In ACM [ACM71], pages 251–257. LCCN QA75.5 .A22 1971.

Maier:1979:MCR

[Mai79] David Maier. Minimum covers in the relational database model (extended abstract). In ACM [ACM79], pages 330–337. LCCN QA75.5.A14 1979.

Manna:1970:SOM

[Man70] Zohar Manna. Second-order mathematical theory of computation. In ACM [ACM70], pages 158–168. LCCN QA75.5 .A22 1970.

Meyer:1972:PSE

[MB72] A. R. Meyer and A. Bagchi. Program size and economy of descriptions: Preliminary Report. In ACM [ACM72], pages 183–186. LCCN QA76.6 .A13 1972.

McGowan:1970:CMS

[McG70] Clement L. McGowan. The correctness of a modified SECD machine. In ACM [ACM70], pages 149–157. LCCN QA75.5 .A22 1970.

McNaughton:1971:DPG

[McN71] Robert McNaughton. A decision procedure for generalized sequential mapability-onto of regular sets. In ACM [ACM71], pages 206–218. LCCN QA75.5 .A22 1971.

Megiddo:1978:COR

[Meg78] Nimrod Megiddo. Combinatorial optimization with rational objective functions. In ACM [ACM78], pages 1–12. LCCN QA76.6.A13 1978.

Mehlhorn:1974:PAS

[Meh74] Kurt Mehlhorn. Polynomial and abstract subrecursive classes. In ACM [ACM74], pages 96–109. LCCN QA76.6 .A13 1974.

Meyers:1971:LRT

[Mey71] W. J. Meyers. Linear representation of tree structure — a mathematical theory of parenthesis-free notations. In ACM [ACM71], pages 50– 62. LCCN QA75.5 .A22 1971.

Miller:1973:TMV

[Mil73] Webb Miller. Toward mechanical verification of properties of roundoff error propagation. In ACM [ACM73], pages 50–58. LCCN QA76.6 .A16 1973.

Millen:1974:CPD

[Mil74a] Jonathan K. Millen. Construction with parallel derivatives of the closure of a parallel program schema. In ACM [ACM74], pages 242–247. LCCN QA76.6 .A13 1974.

Miller:1974:CCN

[Mil74b] Webb Miller. Computational complexity and numerical stability. In ACM [ACM74], pages 317–322. LCCN QA76.6 .A13 1974.

Miller:1975:RHT

[Mil75] Gary L. Miller. Riemann’s Hypothesis and tests for primality. In ACM [ACM75], pages 234–239. LCCN QA76.6 .A16 1975.

Miller:1977:GIG

[Mil77] Gary L. Miller. Graph isomorphism, general remarks. In ACM [ACM77], pages 143–150. LCCN QA76.6 .A13

Miller:1978:ITP

[Mil78] Gary L. Miller. On the n log n isomorphism technique (A preliminary report). In ACM [ACM78], pages 51–58. LCCN QA76.6.A13 1978.

Moenck:1973:FCG

[Moe73] R. T. Moenck. Fast computation of GCDs. In ACM [ACM73], pages 142–151. LCCN QA76.6 .A16 1973.

Morris:1970:RRB

[Mor70] James B. Morris. A result on the relationship between simple precedence languages and reducing transition languages. In ACM [ACM70], pages 73–80. LCCN QA75.5 .A22 1970.

Morris:1972:RSL

[Mor72] James H. Morris. Recursion schemes with lists. In ACM [ACM72], pages 35–43. LCCN QA76.6 .A13 1972.

Manna:1975:OFR

[MS75] Zohar Manna and Adi Shamir. The optimal fixedpoint of recursive programs. In ACM [ACM75], pages 194–206. LCCN QA76.6 .A16 1975.

Munro:1979:IDS

[MS79] J. Ian Munro and Hendra Suwanda.

Implicit data structures (preliminary draft). In ACM [ACM79], pages 108–117. LCCN QA75.5.A14 1979.

Munro:1971:SRC

[Mun71] Ian Munro. Some results concerning efficient and optimal algorithms. In ACM [ACM71], pages 40–44. LCCN QA75.5 .A22 1971.

Martin:1970:SDT

[MV70] David F. Martin and Steven A. Vere. On syntax-directed transduction and tree transducers. In ACM [ACM70], pages 129–135. LCCN QA75.5 .A22 1970.

Meyer:1979:EPD

[MW79] Albert R. Meyer and Karl Winklmann. On the expressive power of Dynamic Logic (preliminary report). In ACM [ACM79], pages 167– 175. LCCN QA75.5.A14 1979.

Machtey:1976:SGN

[MY76] Michael Machtey and Paul Young. Simple Gödel numberings, translations, and the p-hierarchy (preliminary report). In ACM [ACM76], pages 236–243. LCCN QA 76.6 A12 1976.

Miller:1978:FSS

[MY78] Raymond E. Miller and Chee K. Yap. On formulating simultaneity for studying parallelism and synchronization. In ACM [ACM78], pages 105–113. LCCN QA76.6.A13 1978.

Nivat:1970:SFL

[Niv70] Maurice Nivat. On some families of languages related to the Dyck language. In ACM [ACM70], pages 221–225. LCCN QA75.5 .A22 1970.

Nievergelt:1972:BST

[NR72] J. Nievergelt and E. M. Reingold. Binary search trees of bounded balance. In ACM [ACM72], pages 137– 142. LCCN QA76.6 .A13 1972.

Oppen:1975:PAA

[OC75] Derek C. Oppen and Stephen A.

Cook. Proving assertions about programs that manipulate data structures. In ACM [ACM75], pages 107– 116. LCCN QA76.6 .A16 1975.

ODonnell:1977:SRS

[O’D77] Mike O’Donnell. Subtree replacement systems: A unifying theory for recursive equations, LISP, lucid and combinatory logic. In ACM [ACM77], pages 295–305. LCCN QA76.6 .A13 1977.

ODonnell:1979:PLT

[O’D79] Michael O’Donnell. A programming language theorem which is independent of Peano Arithmetic. In ACM [ACM79], pages 176–188. LCCN QA75.5.A14 1979.

Oppen:1973:EBP

[Opp73] Derek C. Oppen. Elementary bounds for Presburger arithmetic. In ACM [ACM73], pages 34–37. LCCN QA76.6 .A16 1973.

Ogden:1972:CTT

[OR72] W. F. Ogden and W. C. Rounds.

Compositions of n tree transducers. In ACM [ACM72], pages 198–206. LCCN QA76.6 .A13 1972.

Owicki:1976:CCD

[Owi76] Susan Owicki. A consistent and complete deductive system for the verification of parallel programs. In ACM [ACM76], pages 73–86. LCCN QA 76.6 A12 1976.

Pager:1973:LTA

[Pag73] David Pager. The lane tracing algorithm for constructing LR(k) parsers. In ACM [ACM73], pages 172–181. LCCN QA76.6 .A16 1973.

Pan:1978:CCC

[Pan78] V. Ya. Pan. Computational complexity of computing polynomials over the fields of real and complex numbers. In ACM [ACM78], pages 162–172. LCCN QA76.6.A13 1978.

Paul:1975:LBC

[Pau75] Wolfgang J. Paul. A 2.5n-lower bound on the combinational complexity of Boolean functions. In ACM [ACM75], pages 27–36. LCCN QA76.6 .A16 1975.

Paul:1977:TH

[Pau77] W. J. Paul. On time hierarchies. In ACM [ACM77], pages 218–222. LCCN QA76.6 .A13 1977.

Perrot:1970:RBF

[Per70] J-F. Perrot. On the relationship between finite automata, finite monoids, and prefix codes. In ACM [ACM70], pages 217–220. LCCN QA75.5 .A22 1970.

Perrault:1975:ITT

[Per75] C. Raymond Perrault. Intercalation theorems for tree transducer languages. In ACM [ACM75], pages 126–136. LCCN QA76.6 .A16 1975.

Peterson:1979:TST

[Pet79] G. L. Peterson. Time-space tradeoffs for asynchronous parallel models (reducibilities and equivalences).

In ACM [ACM79], pages 224–230. LCCN QA75.5.A14 1979.

Peterson:1977:ESC

[PF77] Gary L. Peterson and Michael J.

Fischer. Economical solutions for the critical section problem in a distributed system (extended abstract). In ACM [ACM77], pages 91–97. LCCN QA76.6 .A13 1977.

Pippenger:1976:RMB

[Pip76] Nicholas Pippenger. The realization of monotone Boolean functions (preliminary version). In ACM [ACM76], pages 204–210. LCCN QA 76.6 A12 1976.

Plaisted:1972:FSC

[Pla72] David A. Plaisted. Flowchart schemata with counters. In ACM [ACM72], pages 44–51. LCCN QA76.6 .A13 1972.

Pratt:1973:CPD

[Pra73] Vaughan R. Pratt. Computing permutations with double-ended queues, parallel stacks and parallel queues. In ACM [ACM73], pages 268–277. LCCN QA76.6 .A16 1973.

Pratt:1974:PNT

[Pra74] Vaughan R. Pratt. The power of negative thinking in multiplying Boolean matrices. In ACM [ACM74], pages 80–83. LCCN QA76.6 .A13 1974.

Pratt:1975:FMA

[Pra75] Terrence Pratt. Four models for the analysis and optimization of program control structures. In ACM [ACM75], pages 167–176. LCCN QA76.6 .A16 1975.

Pratt:1978:PDM

[Pra78] V. R. Pratt. A practical decision method for propositional dynamic logic (preliminary report). In ACM [ACM78], pages 326–337. LCCN QA76.6.A13 1978.

Pratt:1974:CPV

[PRS74] Vaughan R. Pratt, Michael O. Rabin, and Larry J. Stockmeyer. A characterization of the power of vector machines. In ACM [ACM74], pages 122–134. LCCN QA76.6 .A13 1974.

Papadimitriou:1976:SCR

[PS76] Christos H. Papadimitriou and Kenneth Steiglitz. Some complexity results for the Traveling Salesman Problem. In ACM [ACM76], pages 1–9. LCCN QA 76.6 A12 1976.

Prabhala:1977:CIS

[PS77] Bhaskaram Prabhala and Ravi Sethi. A comparison of instruction sets for stack machines. In ACM [ACM77], pages 132–142. LCCN QA76.6 .A13 1977.

Paul:1976:SBG

[PTC76] Wolfgang J. Paul, Robert Endre Tarjan, and James R. Celoni. Space bounds for a game on graphs. In ACM [ACM76], pages 149–160. LCCN QA 76.6 A12 1976.

Paterson:1976:LU

[PW76] M. S. Paterson and M. N. Wegman. Linear unification. In ACM [ACM76], pages 181–186. LCCN QA 76.6 A12 1976.

Rackoff:1974:CTW

[Rac74] Charles Rackoff. On the complexity of the theories of weak direct products (preliminary report). In ACM [ACM74], pages 149–160. LCCN QA76.6 .A13 1974.

Rackoff:1978:RQI

[Rac78] Charles Rackoff. Relativized questions involving probabilistic algorithms. In ACM [ACM78], pages 338–342. LCCN QA76.6.A13 1978.

Rajlich:1971:APG

[Raj71] Vaclav Rajlich. Absolutely parallel grammars and two-way deterministic finite-state transducers. In ACM [ACM71], pages 132–137. LCCN QA75.5 .A22 1971.

Reiter:1970:PES

[Rei70] Raymond Reiter. The predicate elimination strategy in theorem proving. In ACM [ACM70], pages 180–183. LCCN QA75.5 .A22 1970.

Reif:1979:UGI

[Rei79] John H. Reif. Universal games of incomplete information. In ACM [ACM79], pages 288–308. LCCN QA75.5.A14 1979.

Reddy:1978:PAB

[RL78] C. R. Reddy and D. W. Loveland. Presburger arithmetic with bounded quantifier alternation. In ACM [ACM78], pages 320–325. LCCN QA76.6.A13 1978.

Rivest:1978:CEB

[RMK78] R. L. Rivest, A. R. Meyer, and D. J. Kleitman. Coping with errors in binary search procedures (preliminary report). In ACM [ACM78], pages 227–232. LCCN QA76.6.A13 1978.

Robertson:1971:CCP

[Rob71] Edward L. Robertson. Complexity classes of partial recursive functions (preliminary version). In ACM [ACM71], pages 258–266. LCCN QA75.5 .A22 1971.

Robertson:1974:SCW

[Rob74] Edward L. Robertson. Structure of complexity in the weak monadic second-order theories of the natural numbers. In ACM [ACM74], pages 161–171. LCCN QA76.6 .A13 1974.

Rosen:1970:TMS

[Ros70a] Barry K. Rosen. Tree-manipulating systems and Church–Rosser Theorems. In ACM [ACM70], pages 117– 127. LCCN QA75.5 .A22 1970.

Rosenberg:1970:DGA

[Ros70b] Arnold L. Rosenberg. Data graphs and addressing schemes (extended abstract). In ACM [ACM70], pages 48–61. LCCN QA75.5 .A22 1970.

Rosenberg:1971:ADG

[Ros71] Arnold L. Rosenberg. Addressable data graphs (extended abstract). In ACM [ACM71], pages 138–150. LCCN QA75.5 .A22 1971.

Rosenberg:1974:MSE

[Ros74] Arnold L. Rosenberg. Managing storage for extendible arrays (extended abstract). In ACM [ACM74], pages 297–302. LCCN QA76.6 .A13 1974.

Rosenthal:1977:NDP

[Ros77] Arnie Rosenthal. Nonserial dynamic programming is optimal. In ACM [ACM77], pages 98–105. LCCN QA76.6 .A13 1977.

Rounds:1970:TOP

[Rou70] William C. Rounds. Tree-oriented proofs of some theorems on contextfree and indexed languages. In ACM [ACM70], pages 109–116. LCCN QA75.5 .A22 1970.

Rosenberg:1975:HSE

[RS75] Arnold L. Rosenberg and Larry J.

Stockmeyer. Hashing schemes for extendible arrays (extended abstract). In ACM [ACM75], pages 159–166. LCCN QA76.6 .A16 1975.

Rose:1975:AAV

[RT75] Donald J. Rose and R. Endre Tarjan. Algorithmic aspects of vertex elimination. In ACM [ACM75], pages 245–254. LCCN QA76.6 .A16 1975.

Ruzzo:1979:TSB

[Ruz79] Walter L. Ruzzo. Tree-size bounded alternation (extended abstract). In ACM [ACM79], pages 352–359. LCCN QA75.5.A14 1979.

Rivest:1975:GPA

[RV75] Ronald L. Rivest and Jean Vuillemin.

A generalization and proof of the Aanderaa-Rosenberg conjecture. In ACM [ACM75], pages 6–11. LCCN QA76.6 .A16 1975.

Raoult:1978:OSE

[RV78] Jean-Claude Raoult and Jean Vuillemin. Operational and semantic equivalence between recursive programs. In ACM [ACM78], pages 75–85. LCCN QA76.6.A13 1978.

Rosenberg:1979:SRT

[RWG79] Arnold L. Rosenberg, Derick Wood, and Zvi Galil. Storage representations for tree-like data structures. In ACM [ACM79], pages 99–107. LCCN QA75.5.A14 1979.

Santos:1971:AST

[San71] Eugene S. Santos. Algebraic structure theory of stochastic machines. In ACM [ACM71], pages 219–243. LCCN QA75.5 .A22 1971.

Savitch:1972:MRA

[Sav72] Walter J. Savitch. Maze recognizing automata (extended abstract). In ACM [ACM72], pages 151–156. LCCN QA76.6 .A13 1972.

Schnorr:1972:PCE

[Sch72] C. P. Schnorr. The process complexity and effective random tests. In ACM [ACM72], pages 168–176. LCCN QA76.6 .A13 1972.

Schaefer:1976:CDP

[Sch76] Thomas J. Schaefer. Complexity of decision problems based on finite two-person perfect-information games. In ACM [ACM76], pages 41– 49. LCCN QA 76.6 A12 1976.

Schaefer:1978:CSP

[Sch78] Thomas J. Schaefer. The complexity of satisfiability problems.

In ACM [ACM78], pages 216–226. LCCN QA76.6.A13 1978.

Seiferas:1974:ONM

[Sei74] Joel I. Seiferas. Observations on nondeterministic multidimensional iterative arrays. In ACM [ACM74], pages 276–289. LCCN QA76.6 .A13 1974.

Sethi:1972:VRA

[Set72] Ravi Sethi. Validating register allocations for straight line programs. In ACM [ACM72], pages 222–237. LCCN QA76.6 .A13 1972.

Sethi:1973:CRA

[Set73] Ravi Sethi. Complete register allocation problems. In ACM [ACM73], pages 182–195. LCCN QA76.6 .A16 1973.

Simon:1977:PRU

[SG77] Istvan Simon and John Gill. Polynomial reducibilities and upward diagonalizations. In ACM [ACM77], pages 186–194. LCCN QA76.6 .A13 1977.

Shamos:1975:GC

[Sha75] Michael Ian Shamos. Geometric complexity. In ACM [ACM75], pages 224–233. LCCN QA76.6 .A16 1975.

Shamir:1979:CKS

[Sha79] Adi Shamir. On the cryptocomplexity of knapsack systems. In ACM [ACM79], pages 118–129. LCCN QA75.5.A14 1979.

Simon:1977:FNP

[Sim77] Janos Simon. On feasible numbers (preliminary version). In ACM [ACM77], pages 195–207. LCCN QA76.6 .A13 1977.

Sipser:1979:LBS

[Sip79] Michael Sipser. Lower bounds on the size of sweeping automata. In ACM [ACM79], pages 360–364. LCCN QA75.5.A14 1979.

Skyum:1975:DLD

[Sky75] Sven Skyum. On decomposing languages defined by parallel devices. In ACM [ACM75], pages 121–125. LCCN QA76.6 .A16 1975.

Stockmeyer:1973:WPR

[SM73] L. J. Stockmeyer and A. R. Meyer. Word problems requiring exponential time (preliminary report). In ACM [ACM73], pages 1–9. LCCN QA76.6 .A16 1973.

Sakoda:1978:NST

[SS78a] William J. Sakoda and Michael Sipser. Nondeterminism and the size of two way finite automata. In ACM [ACM78], pages 275–286. LCCN QA76.6.A13 1978.

Storer:1978:MMD

[SS78b] James A. Storer and Thomas G. Szymanski. The macro model for data compression (extended abstract). In ACM [ACM78], pages 30–39. LCCN QA76.6.A13 1978.

Sedgewick:1979:CFP

[SS79] Robert Sedgewick and Thomas G. Szymanski. The complexity of finding periods. In ACM [ACM79], pages 74–80. LCCN QA75.5.A14

Sacerdote:1977:DRP

[ST77] George S. Sacerdote and Richard L.

Tenney. The decidability of the reachability problem for vector addition systems (preliminary version). In ACM [ACM77], pages 61– 76. LCCN QA76.6 .A13 1977.

Sciore:1978:CTA

[ST78] E. Sciore and A. Tang. Computability theory in admissible domains. In ACM [ACM78], pages 95–104. LCCN QA76.6.A13 1978.

Stanat:1971:FLP

[Sta71] D. F. Stanat. Formal languages and power series. In ACM [ACM71], pages 1–11. LCCN QA75.5 .A22 1971.

Strong:1970:TRE

[Str70] H. R. Strong. Translating recursion equations into flow charts. In ACM [ACM70], pages 184–197. LCCN QA75.5 .A22 1970.

Sudborough:1976:DCF

[Sud76] I. H. Sudborough. On deterministic context-free languages, multihead automata, and the power of an auxiliary pushdown store. In ACM [ACM76], pages 141–148. LCCN QA 76.6 A12 1976.

Sudborough:1977:STB

[Sud77] I. H. Sudborough. Separating tape bounded auxiliary pushdown automata classes. In ACM [ACM77], pages 208–217. LCCN QA76.6 .A13

Symes:1972:CFF

[Sym72] D. M. Symes. The computation of finite functions. In ACM [ACM72], pages 177–182. LCCN QA76.6 .A13 1972.

Tarjan:1973:TFG

[Tar73] Robert Tarjan. Testing flow graph reducibility. In ACM [ACM73], pages 96–107. LCCN QA76.6 .A16 1973.

Tarjan:1974:TGC

[Tar74] R. Endre Tarjan. Testing graph connectivity. In ACM [ACM74], pages 185–193. LCCN QA76.6 .A13 1974.

Tarjan:1977:RMR

[Tar77] Robert Endre Tarjan. Reference machines require non-linear time to maintain disjoint sets. In ACM [ACM77], pages 18–29. LCCN QA76.6 .A13 1977.

Teitelbaum:1973:CFE

[Tei73] Ray Teitelbaum. Context-free error analysis by evaluation of algebraic power series. In ACM [ACM73], pages 196–199. LCCN QA76.6 .A16 1973.

Terrine:1971:AGD

[Ter71] Gérard Terrine. An algorithm generating the decision table of a deterministic bottom up parser for a subset of context free grammars. In ACM [ACM71], pages 185–205. LCCN QA75.5 .A22 1971.

Thompson:1979:ATC

[Tho79] C. D. Thompson. Area-time complexity for VLSI. In ACM [ACM79], pages 81–88. LCCN QA75.5.A14 1979.

Thompson:1976:SMC

[TK76] C. D. Thompson and H. T. Kung.

Sorting on a mesh-connected parallel computer. In ACM [ACM76], pages 58–64. LCCN QA 76.6 A12 1976.

Tompa:1978:TST

[Tom78] Martin Tompa. Time-space tradeoffs for computing functions, using connectivity properties of their circuits. In ACM [ACM78], pages 196– 204. LCCN QA76.6.A13 1978.

Toueg:1979:DFP

[TU79] Sam Toueg and Jeffrey D. Ullman.

Deadlock-free packet switching networks. In ACM [ACM79], pages 89– 98. LCCN QA75.5.A14 1979.

Thatcher:1978:DTS

[TWW78] James W. Thatcher, Eric G. Wagner, and Jesse B. Wright. Data type specification: Parameterization and the power of specification techniques. In ACM [ACM78], pages 119–132. LCCN QA76.6.A13 1978.

Ullian:1970:TTA

[Ull70] Joseph S. Ullian. Three theorems on abstract families of languages. In ACM [ACM70], pages 226–230. LCCN QA75.5 .A22 1970.

Valiant:1974:DED

[Val74] Leslie G. Valiant. The decidability of equivalence for deterministic finite-turn pushdown automata. In ACM [ACM74], pages 27–32. LCCN QA76.6 .A13 1974.

Valiant:1975:NLL

[Val75] Leslie G. Valiant. On non-linear lower bounds in computational complexity. In ACM [ACM75], pages 45–53. LCCN QA76.6 .A16 1975.

Valiant:1976:UCP

[Val76] Leslie G. Valiant. Universal circuits (preliminary report). In ACM [ACM76], pages 196–203. LCCN QA 76.6 A12 1976.

Valiant:1979:CCA

[Val79a] L. G. Valiant. Completeness classes in algebra. In ACM [ACM79], pages 249–261. LCCN QA75.5.A14 1979.

Valiant:1979:NCE

[Val79b] L. G. Valiant. Negation can be exponentially powerful. In ACM [ACM79], pages 189–196. LCCN QA75.5.A14 1979.

Vairavan:1974:PSP

[VD74] K. Vairavan and R. A. DeMillo. Parallel scheduling of programs in a restricted model of computation. In ACM [ACM74], pages 248–255. LCCN QA76.6 .A13 1974.

vanLeeuwen:1974:PSR

[vL74] Jan van Leeuwen. A partial solution to the reachability-problem for vector-addition systems. In ACM [ACM74], pages 303–309. LCCN QA76.6 .A13 1974.

Valdes:1979:RSP

[VTL79] Jacobo Valdes, Robert E. Tarjan, and Eugene L. Lawler. The recognition of Series Parallel digraphs. In ACM [ACM79], pages 1–12. LCCN QA75.5.A14 1979.

Vuillemin:1973:COI

[Vui73] Jean Vuillemin. Correct and optimal implementations of recursion in a simple programming language. In ACM [ACM73], pages 224–239. LCCN QA76.6 .A16 1973.

Wagner:1971:ATR

[Wag71] Eric G. Wagner. An algebraic theory of recursive definitions and recursive languages. In ACM [ACM71], pages 12–23. LCCN QA75.5 .A22 1971.

Wagner:1973:APL

[Wag73] Eric G. Wagner. From algebras to programming languages. In ACM [ACM73], pages 214–223. LCCN QA76.6 .A16 1973.

Wagner:1975:CES

[Wag75] Robert A. Wagner. On the complexity of the Extended String-to-String Correction Problem. In ACM [ACM75], pages 218–223. LCCN QA76.6 .A16 1975.

Wand:1973:UAP

[Wan73] Mitchell Wand. An unusual application of program-proving. In ACM [ACM73], pages 59–66. LCCN QA76.6 .A16 1973.

Wand:1976:NIR

[Wan76] Mitchell Wand. A new incompleteness result for Hoare’s system. In ACM [ACM76], pages 87–91. LCCN QA 76.6 A12 1976.

Wegener:1978:SFW

[Weg78] Ingo Wegener. Switching functions whose monotone complexity.

In ACM [ACM78], pages 143–149. LCCN QA76.6.A13 1978.

Warkentin:1972:PMR

[WF72] John C. Warkentin and Patrick C. Fischer. Predecessor machines and regressing functions. In ACM [ACM72], pages 81–87. LCCN QA76.6 .A13 1972.

Wise:1971:DAA

[Wis71] David S. Wise. Domolki’s algorithm applied to generalized overlap resolvable grammars. In ACM [ACM71], pages 171–184. LCCN QA75.5 .A22 1971.

Wotschke:1975:DLP

[Wot75] Detlef Wotschke. Degree-languages, polynomial time recognition, and the LBA problem. In ACM [ACM75], pages 145–152. LCCN QA76.6 .A16 1975.

Walker:1972:CFR

[WS72] S. A. Walker and H. R. Strong. Characterizations of flowchartable recursions short version. In ACM [ACM72], pages 18–34. LCCN QA76.6 .A13 1972.

Weiner:1973:NSA

[WSB73] P. Weiner, S. L. Savage, and A. Bagchi. Neighborhood search algorithms for finding optimal traveling salesman tours must be inefficient. In ACM [ACM73], pages 207– 213. LCCN QA76.6 .A16 1973.

Yannakakis:1978:NED

[Yan78] Mihalis Yannakakis. Node-and edge-deletion NP-complete problems. In ACM [ACM78], pages 253– 264. LCCN QA76.6.A13 1978.

Yao:1975:CMQ

[Yao75] Andrew Chi-Chih Yao. On computing the minima of quadratic forms (preliminary report). In ACM [ACM75], pages 23–26. LCCN QA76.6 .A16 1975.

Yao:1979:SCQ

[Yao79] Andrew Chi-Chih Yao. Some complexity questions related to distributive computing (preliminary report). In ACM [ACM79], pages 209–213. LCCN QA75.5.A14 1979.

Yao:1977:LBS

[YAR77] Andrew C. Yao, David M. Avis, and Ronald L. Rivest. An Ω(n^{2} log n) lower bound to the shortest paths problem. In ACM [ACM77], pages 11–17. LCCN QA76.6 .A13 1977.